# Integrating Various Software Artifacts for Better LLM-based Bug Localization and Program Repair

### v2

@@ Abstract.
To investigate this and mimic the way human developers fix bugs, we propose DEVLoRe (short for DEVeloper Localization and Repair). In this framework, LLMs first use issue content (description and discussion) and stack error traces to localize buggy methods, then rely on debug information in buggy methods and issue content and stack error to localize buggy lines and generate valid patches. We evaluated the effectiveness of issue content, error stack traces, and debugging information in bug localization and automatic program repair. Our results show that while issue content and error stack is particularly effective in assisting LLMs with fault localization and program repair respectively, different types of software artifacts complement each other in addressing various bugs. By incorporating these three types of artifacts and using the Defects4J v2.0 dataset for evaluation, DEVLoRe successfully localizes 49.3% of single-method bugs and generates 56.0% plausible patches. Additionally, DEVLoRe can localize 47.6% of non-single-method bugs and generates 14.5% plausible patches. Moreover, our framework streamlines the end-to-end process from buggy source code to a complete repair, and achieves a 39.7% and 17.1% of single-method and non-single-method bug repair rate, outperforming current state-of-the-art APR methods. The source code and experimental results of this work for replication are available at https://github.com/XYZboom/DEVLoRe.
@@
copyright: acmlicensedjournal: TOSEMjournalyear: 2024doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Software verification and validation
@@ 1. Introduction
Recent studies have explored the use of LLMs for APR, either by having LLMs fill in the correct code in buggy methods using an infilling-style technique or by directly generating patches when provided with buggy methods (Xia and Zhang, 2023; Xia et al., 2024a, b; Li et al., 2024; Xia et al., 2023b, a; Jiang et al., 2023; Wei et al., 2023; Zhang et al., 2023; Silva et al., 2023). The initial results demonstrate the ability of LLMs to correctly repair real-world bugs, including those that were previously unrepairable by existing APR approaches. One of the best-performing frameworks is Agentless (Xia et al., 2024a), which feeds an LLM with only issue description and can automatically solve GitHub issues. These promising outcomes highlight the potential of LLMs to develop more effective APR methods.
@@
    To the best of our knowledge, this work is the first to compare different software artifacts in assisting LLMs’ ability to perform fault localization and program repair. The results can help developers understand LLMs’ potential when provided with a variety of software artifacts.
@@
    We propose a simple and lightweight framework that leverages LLMs to conduct an end-to-end process for bug localization and program repair. Accompanying this framework we design a strict input/output prompt. The DEVLoRe framework demonstrates a strong ability to localize and fix more software bugs in less time and at a lower cost, compared to current state-of-the-art methods.
@@ 3.4. Patch Validation
To be consistent with existing studies (Zhu et al., 2023; Li et al., 2024), if the patch is semantically equivalent to the original patch provided by developers, it is considered a correct patch. As part of this validation process, the patch undergoes manual inspection and cross-checking by two experienced developers to ensure its correctness. This step is crucial to verify that the patch not only passes all automated tests but also aligns with the intended behavior from a developer’s perspective. The manual inspection by developers serves as a final quality control step to ensure that the patch addresses the bug correctly and does not introduce new bugs.
@@ 4.1. Dataset
4.2. Parameters in Experiments
@@ 5.1.1. Localizing Buggy Method(s)
Table 2 shows the effectiveness of buggy method localization in these four experiments. The first row, labeled “—”, indicates that only the related (executed) methods are fed to the LLM for buggy method localization. The second, third and fourth rows represent scenarios when, in addition to method localization, issue content, error stack, or both are provided. Following the approach described in LLMAO (Yang et al., 2024a), we consider a buggy method correctly localized if at least one match in the LLM’s response corresponds to the buggy method (single-method bugs) or one of the buggy methods (non-single-method bugs) in the ground truth. It is worth mentioning that there exist bug cases where certain software artifacts can not be extracted, and in this case we calculate the ratio by dividing only the number of bug cases with certain software artifacts information. Table 2 shows that, when issue content or error stack is provided, the LLM is able to localize more buggy methods than nothing is provided. When issue content is provided, 15.8% more single buggy methods and 16.5% more non-single buggy methods are correctly localized. When error stack is included, 16.6% more single buggy methods and 14.3% more non-single buggy methods are correctly localized. Furthermore, combining issue content and error stack achieves the best performance, which results in 49.3% in localizing single buggy methods and 47.6% of non-single buggy methods.
Table 3. Method-level FL comparision with baselines on Defects4J v1.2
Projects #Bugs DEVLoRe AgentFL (Qin et al., 2024) GRACE (Lou et al., 2021) FLUCCS (Sohn and Yoo, 2017)
@@
Chart 26 10 18 19 16 18 19 14 20 22 15 16 19
Lang 65 47 52 53 44 45 45 42 54 57 40 53 55
Math 106 71 83 84 49 60 61 61 78 89 48 77 83
Time 27 11 14 15 11 13 13 11 14 19 8 15 18
Mockito 38 16 21 22 13 14 14 17 24 26 7 19 22
Closure 133 19 32 35 24 33 35 47 70 81 42 66 77
Total 395 174 220 228 157 183 187 192 260 294 160 249 271
w.o Closure 262 155 188 193 133 150 152 145 190 213 118 183 194
@@
As existing SOTA fault localization approaches use Defects4J v1.2 dataset and adopt Top-n (the buggy method is in the top $n$ list of the LLM’s response) to evaluate the performance of fault localization, we compare the performance of DEVLoRe with the SOTA approaches in the same dataset and adopt Top-n approach. Table 3 shows the comparison results where AgentFL (Qin et al., 2024) is LLM-based, GRACE (Lou et al., 2021) and FLUCCS (Sohn and Yoo, 2017) are learning-based. As we can see,  DEVLoRe outperforms another LLM-based approach AgentFL (Qin et al., 2024) in Top-1, Top-3 and Top-5 metrics. As discussed in (Qin et al., 2024), the Closure project contains many bugs with similar code, making it more suitable for learning-based approaches. This characteristic allows GRACE (Lou et al., 2021) and FLUCCS (Sohn and Yoo, 2017) to achieve an overall high localization rate. The last row shows that, without the Closure project, DEVLoRe outperforms other approaches in Top-1 and achieves similar performance in Top-3 and Top-5, demonstrating its ability to precisely localize buggy methods.
@@ 5.1.2. Localizing Buggy Line(s)
Table 4 shows the effectiveness of buggy line localization with different artifacts. We can see that when the LLM is fed with issue content, it performs the best among all single software artifacts, achieving 49.5% in exact matches, 74.4% in Range-3, and 77.5% in Range-5 for single methods. When adding more software artifacts, the performance can be further increased. For example, when adding issue content to error stack, the exact, Range-3, and Range-5 buggy line match for single methods can be increased from 45.0% to 49.1%, 66.9% to 73.2%, and 71.0% to 77.5%. The last row indicates that, when combining all artifacts, 68.9% of buggy lines in single-method bugs and 9% of buggy lines in non-single-method bugs are correctly localized. Similar to Figure 7, Figure 8 demonstrates that, while most of the correctly localized buggy lines across different artifacts overlap, different combinations of artifacts each have their own advantages (since only five areas can be displayed in the overlap figure, we randomly select five categories for visualization). For example, using only the issue content can correctly localize 7 more buggy lines, while using only the debug information can localize 15 more.
@@
Table 6 compares DEVLoRe with other SOTA program repair approaches. As these approaches focus on repairing single-method bugs, we compare our tool with these approaches on the same dataset. As shown in Table 6, 142 out of 259 bugs in Defects4J v1.2 and 132 out of 230 bugs in Defects4J v2.0 can be repaired by DEVLoRe, outperforming the other five approaches. In total, DEVLoRe can fix 274 buggy methods, which is 60.2% more than GiantRepair (Li et al., 2024). Furthermore, Figure 10 shows that 105 bugs can only be fixed by DEVLoRe. This result demonstrates DEVLoRe’s ability in fixing more and extra bugs by leveraging different types of software artifacts.
Table 6. Plausible patch with provided method-level fault localization in single buggy method projects.
Project #Bugs DEVLoRe GaintRepair (Li et al., 2024) FitRepair (Xia et al., 2023a) Repilot (Wei et al., 2023) Tare (Zhu et al., 2023) GAMMA (Zhang et al., 2023)
Chart 16 12 8 8 6 11 9
Closure 95 37 32 29 21 22 20
Lang 42 32 14 17 15 13 10
Math 74 53 26 23 20 20 19
Time 16 3 1 3 2 3 1
Mockito 16 10 6 4 0 2 2
Defects4J v1.2 259 142 87 85 64 71 61
Defects4J v2.0 230 132 84 44 47 37 39
Total 489 274 171 129 111 108 100
Figure 10. A Comparison between DEVLoRe with Other State-of-the-art SOTA Program Repair Approaches
@@ 6.1.1. A simple but efficient framework:
Unlike approaches that rely on patch skeletons (Li et al., 2024), fix templates (Zhang et al., 2023), or type checking (Zhu et al., 2023), DEVLoRe adopts a straightforward framework that allows LLMs to handle both fault localization and program repair with the aid of two lightweight tools we implemented MethodRecorder and DebugRecorder, making it both simple and efficient. The findings of RQ1, RQ2, and RQ3 demonstrate that different software artifacts may lead to different performance in bug localization or program repair. Also, because there is no constraints such as fill-in-the-blank templates or code skeletons for generating patches (Li et al., 2024; Zhang et al., 2023), DEVLoRe can fix bugs that other tools cannot, especially some non-single method bugs. Our results in Table 7 show that by combining issue content, stack error, and debug information, our approach can fix 11.2% of non-single method bugs. More importantly, combinations of software artifacts can achieve the best performance in all experiments: bug localization, program repair, and the overall streamlined process. One explanation for the strong performance of DEVLoRe is that by integrating multiple sources of information, the noise inherent in any single source can be significantly reduced (Meng et al., 2024). This allows the LLM to focus on the consistent and complementary aspects of the different artifacts, enabling it to perform more effective reasoning and making better use of the available data. While it may be argued that the good results of DEVLoRe are due to the use of GPT-4 models, a recent study shows that directly using the GPT-4 model does not improve the fix rate (Li et al., 2024). Therefore, we argue that our proposed framework is the key to achieving the high rate in fault localization and program repair.
@@ 6.1.2. A strict input/output prompt design:
Project DEVLoRe GaintRepair (Li et al., 2024) Tare (Zhu et al., 2023) TBar (Liu et al., 2019)
@@ 6.1.3. Plausible patch to correct patch:
Most state-of-the-art (SOTA) program repair approaches use plausible patches that pass all unit tests as an important evaluation metric, since generating plausible patches within limited time and resources is crucial for practical applications. To facilitate a better comparison, we also used plausible patches in RQ2 (Which software artifacts can better assist LLMs in generating valid patches when provided with buggy methods?) and RQ3 (What is the overall performance of DEVLoRe in bug localization and program repair?). Some may argue that a high number of plausible patches does not necessarily translate to a high number of semantically correct patches. To address this, we manually inspected the plausible patches, using the approach described in Section 3.4. Table 9 compares the end-to-end repair results on Defects4J v1.2, as all the baselines consistently used this benchmark. Our DEVLoRe framework uses the LLM to localize buggy methods and lines, whereas all three baselines — GiantRepair (Li et al., 2024), Tare (Zhu et al., 2023), and TBar (Liu et al., 2019) — employ the spectrum-based algorithm Ochiai (Abreu et al., 2007), implemented by GZoltar (Campos et al., 2012), to localize buggy methods. As shown in Table 9, the ratio of correct to plausible patches (56.2%) produced by our DEVLoRe is lower than that from Tare (60.0%) and from GiantRepair (57.7%). This difference is understandable, as GiantRepair enforces an AST patch skeleton and Tare uses a typing-checking mechanism. Notably, 59 out of 105 plausible patches generated by DEVLoRe are evaluated as correct patches, which is higher than the 57 in Tare and 31 in TBar. This finding demonstrates DEVLoRe’s ability to generate candidate patches that not only pass all unit tests but are also semantically correct.

### v3

@@ Abstract.
To investigate this and mimic the way human developers fix bugs, we propose DEVLoRe (short for DEVeloper Localization and Repair). In this framework, LLMs first use issue content (description and discussion) and stack error traces to localize buggy methods, then rely on debug information in buggy methods and issue content and stack error to localize buggy lines and generate valid patches. We evaluated the effectiveness of issue content, error stack traces, and debugging information in bug localization and automatic program repair. Our results show that while issue content and error stack is particularly effective in assisting LLMs with fault localization and program repair respectively, different types of software artifacts complement each other in addressing various bugs. By incorporating these three types of artifacts and using the Defects4J v2.0 dataset for evaluation, DEVLoRe successfully localizes 49.3% of single-method bugs and generates 56.0% plausible patches. Additionally, DEVLoRe can localize 47.6% of non-single-method bugs and generates 14.5% plausible patches. Moreover, our framework streamlines the end-to-end process from buggy source code to a complete repair, and achieves a 39.7% and 17.1% of single-method and non-single-method bug repair rate, outperforming current state-of-the-art APR methods. Furthermore, we re-implemented and evaluated our framework, demonstrating its effectiveness in its effectiveness in resolving 9 unique issues compared to other state-of-the-art frameworks using the same or more advanced models on SWE-bench Lite. We also discussed whether a leading framework for Python code can be directly applied to Java code, or vice versa. The source code and experimental results of this work for replication are available at https://github.com/XYZboom/DEVLoRe.
@@
copyright: acmlicensedjournal: TOSEMjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Software verification and validation
@@ 1. Introduction
Recent studies have explored the use of LLMs for APR, either by having LLMs fill in the correct code in buggy methods using an infilling-style technique or by directly generating patches when provided with buggy methods (Xia and Zhang, 2023; Xia et al., 2024a, b; Li et al., 2025; Xia et al., 2023b, a; Jiang et al., 2023; Wei et al., 2023; Zhang et al., 2023; Silva et al., 2023). The initial results demonstrate the ability of LLMs to correctly repair real-world bugs, including those that were previously unrepairable by existing APR approaches. One of the best-performing frameworks is Agentless (Xia et al., 2024a), which feeds an LLM with only issue description and can automatically solve GitHub issues. These promising outcomes highlight the potential of LLMs to develop more effective APR methods.
@@
    To the best of our knowledge, this work is the first to compare different software artifacts in assisting LLMs’ ability to perform both fault localization and program repair. The results can help developers understand LLMs’ potential when provided with a variety of software artifacts.
@@
    We propose a simple and lightweight framework that leverages LLMs to conduct an end-to-end process for bug localization and program repair. Accompanying this framework we design a strict input/output prompt. The DEVLoRe framework demonstrates a strong ability to localize and fix more software bugs in less time and at a lower cost, compared to current state-of-the-art methods on the Defects4J dataset.
  3. (3)
    We evaluated our framework on SWE-bench Lite and demonstrated DEVLoRe’s effectiveness in resolving 9 unique issues compared to other state-of-the-art frameworks using the same or more advanced models. We further discussed whether a leading framework for Python code can be directly applied to Java code, or vice versa..
@@ 3.4. Patch Validation
To be consistent with existing studies (Zhu et al., 2023; Li et al., 2025), if the patch is semantically equivalent to the original patch provided by developers, it is considered a correct patch. As part of this validation process, the patch undergoes manual inspection and cross-checking by two experienced developers to ensure its correctness. This step is crucial to verify that the patch not only passes all automated tests but also aligns with the intended behavior from a developer’s perspective. The manual inspection by developers serves as a final quality control step to ensure that the patch addresses the bug correctly and does not introduce new bugs.
@@ 4.1. Dataset
4.2. Parameters in Experiments and Baseline Selection
@@ 5.1.1. Localizing Buggy Method(s)
Table 2 shows the effectiveness of buggy method localization in these four experiments. The first row, labeled “—”, indicates that only the related (executed) methods are fed to the LLM for buggy method localization. The second, third and fourth rows represent scenarios when, in addition to related method localization, issue content, error stack, or both are provided. Following the approach described in LLMAO (Yang et al., 2024b), we consider a buggy method correctly localized if at least one match in the LLM’s response corresponds to the buggy method (single-method bugs) or one of the buggy methods (non-single-method bugs) in the ground truth. It is worth mentioning that there exist bug cases where certain software artifacts can not be extracted, and in this case we calculate the ratio by dividing only the number of bug cases with certain software artifacts information. Table 2 shows that, when issue content or error stack is provided, the LLM is able to localize more buggy methods than nothing is provided. When issue content is provided, 15.8% more single buggy methods and 16.5% more non-single buggy methods are correctly localized. When error stack is included, 16.6% more single buggy methods and 14.3% more non-single buggy methods are correctly localized. Furthermore, combining issue content and error stack achieves the best performance, which results in 49.3% in localizing single buggy methods and 47.6% of non-single buggy methods.
Table 3. Method-level FL comparison with SOTA baselines on Defects4J v1.2
Project #Bugs DEVLoRe* SoapFL (Qin et al., 2025) AutoFL (Kang et al., 2024) Agentless (Xia et al., 2024a)
codebase, issue content, failed tests, error stack codebase, failed tests, error stack failed tests, error stack codebase, issue message
@@
Chart 26 18 18 18 17 22 22 16 20 21 10 15 16
Lang 65 54 58 58 39 48 49 35 43 43 34 40 48
Math 106 81 86 86 57 71 73 53 67 70 45 51 52
Time 27 12 13 13 10 13 13 14 15 16 8 10 12
Mockito 38 22 22 22 20 21 22 13 17 17 21 26 26
Closure 133 26 37 41 32 57 57 18 24 24 18 31 31
Overall 395 213 234 238 175 232 236 149 186 191 136 173 185
*: To ensure a fair comparison, DEVLoRe uses the same GPT-3.5 model as the other baselines.
@@
As existing SOTA fault localization approaches use Defects4J v1.2 dataset and adopt Top-n (the buggy method is in the top $n$ list of the LLM’s response) to evaluate the performance of fault localization, we compare the performance of DEVLoRe with the SOTA approaches in the same dataset and adopt Top-n approach. Table 3 presents the comparison results of DEVLoRe with three state-of-the-art (SOTA) LLM-based approaches. The row below the names of these approaches lists the artifacts used by each approach. Among them, SoapFL (Qin et al., 2025) uses codebase, failed test methods, and error stack trace; AutoFL (Kang et al., 2024) uses failed test methods along with error stack trace; and Agentless (Xia et al., 2024a) uses codebase and issue message. In contrast to these approaches, DEVLoRe utilizes four artifacts: issue content, error stack, failed tests, and codebase. Failed tests and codebase are used to retrieve related methods via MethodRecorder. It is worth to mention that DEVLoRe also outperform the SOTA learning-based approaches such as GRACE (Lou et al., 2021) and FLUCCS (Sohn and Yoo, 2017). But in this study we focused on how various artifacts can help LLMs with bug localization and program repair, so we did not include the comparison with non-LLM-based approaches. Since the other three baselines use the GPT-3.5 model, for a fair comparison, DEVLoRe also uses the same GPT-3.5 model. As we can see, DEVLoRe outperforms SoapFL, AutoFL and Agentless in Top-1, Top-3 and Top-5 metrics. The last row shows that, DEVLoRe outperforms other approaches in Top-1 and achieves similar performance in Top-3 and Top-5 with SoapFL, demonstrating its ability to precisely localize buggy methods.
@@ 5.1.2. Localizing Buggy Line(s)
Table 4 shows the effectiveness of buggy line localization with different artifacts. We can see that when the LLM is fed with issue content, it performs the best among all single software artifacts, achieving 49.5% in exact matches, 74.4% in Range-3, and 77.5% in Range-5 for single methods. When adding more software artifacts, the performance does not appear to improve. For example, combining issue content with the error stack does not yield better exact, Range-3, or Range-5 buggy line matches for single methods compared to using issue content alone. However, as seen from Figure 8, the combination does successfully localize 3 additional buggy lines that cannot be identified using either issue content or the error stack individually. The last row indicates that, when combining all artifacts, 68.9% of buggy lines in single-method bugs and 9% of buggy lines in non-single-method bugs are correctly localized. Similar to Figure 7, Figure 8 demonstrates that, while most of the correctly localized buggy lines across different artifacts overlap, different combinations of artifacts each have their own advantages (since only five areas can be displayed in the overlap figure, we randomly select five categories for visualization). For example, using only the issue content can correctly localize 7 more buggy lines, while using only the debug information can localize 15 more.
@@
Table 6 compares DEVLoRe with other state-of-the-art (SOTA) LLM-based program repair approaches. Since these approaches focus on repairing single-method bugs, we evaluate our tool on the same dataset for a fair comparison. The row below the names of these approaches lists the artifacts used by each approach. Specifically, GiantRepair (Li et al., 2025) and Repilot (Wei et al., 2023) take the buggy function as input; RepairAgent (Bouzenia et al., 2024) uses the failed tests and all methods in the codebase; ChatRepair (Xia and Zhang, 2023) uses the failed tests and the buggy line; and FitRepair (Xia et al., 2023a) uses only the buggy lines within the buggy method. In contrast, DEVLoRe leverages four artifacts including the buggy method, debugging information, stack traces, and issue content. As shown in Table 6, 142 out of 259 bugs in Defects4J v1.2 and 132 out of 230 bugs in Defects4J v2.0 can be repaired by DEVLoRe, outperforming the other five LLM-based approaches. In total, DEVLoRe can fix 274 buggy methods, which is 60.2% more than GiantRepair (Li et al., 2025). Furthermore, Figure 10 shows that 88 bugs can only be fixed by DEVLoRe. This result demonstrates DEVLoRe’s ability in fixing more and extra bugs by leveraging different types of software artifacts.
Table 6. Plausible patches with provided method-level fault localization in single buggy method projects
Project #Bugs DEVLoRe GaintRepair RepairAgent ChatRepair FitRepair Repilot
buggy method, issue content, debugging info, stack traces buggy method failed tests, all methods in codebase failed tests, buggy lines buggy lines within buggy method buggy method
Chart 16 12 8 11 15 8 6
Closure 95 37 32 27 37 29 21
Lang 42 32 14 17 21 17 15
Math 74 53 26 29 32 23 20
Time 16 3 1 2 6 3 2
Mockito 16 10 6 6 3 4 0
Defects4J v1.2 259 142 87 74 114 85 64
Defects4J v2.0 230 132 84 90 48 44 47
Total 489 274 171 164 162 129 111
Figure 10. A Comparison between DEVLoRe with other LLM-based SOTA program repair approaches
@@ 6.1.1. A simple but efficient framework:
Unlike approaches that rely on patch skeletons (Li et al., 2025), fix templates (Zhang et al., 2023), or type checking (Zhu et al., 2023), DEVLoRe adopts a straightforward framework that allows LLMs to handle both fault localization and program repair with the aid of two lightweight tools we implemented MethodRecorder and DebugRecorder, making it both simple and efficient. The findings of RQ1, RQ2, and RQ3 demonstrate that different software artifacts may lead to different performance in bug localization or program repair. Also, because there is no constraints such as fill-in-the-blank templates or code skeletons for generating patches (Li et al., 2025; Zhang et al., 2023), DEVLoRe can fix bugs that other tools cannot, especially some non-single method bugs. Our results in Table 7 show that by combining issue content, stack error, and debug information, our approach can fix 11.2% of non-single method bugs. More importantly, combinations of software artifacts can achieve the best performance in all experiments: bug localization, program repair, and the overall streamlined process. One explanation for the strong performance of DEVLoRe is that by integrating multiple sources of information, the noise inherent in any single source can be significantly reduced (Meng et al., 2024). This allows the LLM to focus on the consistent and complementary aspects of the different artifacts, enabling it to perform more effective reasoning and making better use of the available data. While it may be argued that the good results of DEVLoRe are due to the use of GPT-4o models, a recent study shows that directly using the GPT-4o model does not improve the fix rate (Li et al., 2025). Therefore, we argue that our proposed framework is the key to achieving the high rate in fault localization and program repair.
@@ 6.1.2. A strict input/output prompt design:
Project DEVLoRe GaintRepair (Li et al., 2025) Tare (Zhu et al., 2023) TBar (Liu et al., 2019)
@@ 6.1.3. Plausible patch to correct patch:
Most state-of-the-art (SOTA) program repair approaches use plausible patches that pass all unit tests as an important evaluation metric, since generating plausible patches within limited time and resources is crucial for practical applications. To facilitate a better comparison, we also used plausible patches in RQ2 (Which software artifacts can better assist LLMs in generating valid patches when provided with buggy methods?) and RQ3 (What is the overall performance of DEVLoRe in bug localization and program repair?). Some may argue that a high number of plausible patches does not necessarily translate to a high number of semantically correct patches. To address this, we manually inspected the plausible patches, using the approach described in Section 3.4. Table 9 compares the end-to-end repair results on Defects4J v1.2, as all the baselines consistently used this benchmark. Our DEVLoRe framework uses the LLM to localize buggy methods and lines, whereas all three approaches — GiantRepair (Li et al., 2025), Tare (Zhu et al., 2023), and TBar (Liu et al., 2019) — employ the spectrum-based algorithm Ochiai (Abreu et al., 2007), implemented by GZoltar (Campos et al., 2012), to localize buggy methods. As shown in Table 9, the ratio of correct to plausible patches (56.2%) produced by our DEVLoRe is lower than that from Tare (60.0%) and from GiantRepair (57.7%). This difference is understandable, as GiantRepair enforces an AST patch skeleton and Tare uses a typing-checking mechanism. Notably, 59 out of 105 plausible patches generated by DEVLoRe are evaluated as correct patches, which is higher than the 57 in Tare and 31 in TBar. This finding demonstrates DEVLoRe’s ability to generate candidate patches that not only pass all unit tests but are also semantically correct.
6.2. Comparison with Baselines of Different Models or on Different Languages
6.2.1. Comparison with Baselines of Different Models
We selected state-of-the-art LLM-based fault localization and program repair frameworks on the Defects4J dataset as our baselines. One may argue that some baselines do not use a GPT-4o model, and that the superior performance of our DEVLoRe framework is due to the use of a more powerful model. However, as shown in Table 3, even using the same GPT-3.5 model as the baselines, our framework outperforms existing state-of-the-art fault localization approaches. For program repair, the models used in these baselines vary significantly. For example, GiantRepair and Repilot combine the outputs of multiple LLMs, while FitRepair fine-tunes CodeT5 using 50% of the Defects4J dataset. As a result, it is challenging to align all baselines under a single model. According to a recent study (Campos et al., 2025), GPT-4o achieves only 16.4% and 27.8% accuracy at pass@1 and pass@5, respectively, in generating correct patches given a buggy function on the Defects4J dataset. Additionally, as noted in the discussion section of the GiantRepair paper (Li et al., 2025), given a buggy function, GPT-4o-mini achieves 21.7% accuracy when generating 200 patches per bug on the Defects4J dataset (pass@200). In contrast, DEVLoRe, which leverages issue content, error stack and debug information, achieves 56.0% accuracy in program repair using GPT-4o-mini (pass@30). We believe that this performance gap is due to our framework’s ability to effectively leverage various software artifacts.
6.2.2. Comparison with Baselines on Different Programming Languages
Our framework was largely inspired by Agentless. Initially, we attempted to apply Agentless to the Defects4J dataset, but the resolution rate was below 10%. This led us to consider incorporating debugging information and re-implementing Agentless for evaluation on the Defects4J dataset. During this process, we encountered a major issue: Agentless uses only the skeleton of the source files during the fault localization step, without incorporating the detailed source code. To address this limitation and incorporate both debugging information and actual code lines, we extended the original Agentless framework to include the full source code as input to the LLM. However, this introduced a new challenge: the input frequently exceeded the LLM’s token limit. To overcome this, we introduced MethodRecorder, which identifies and retains only the methods relevant to the failing test cases, thereby reducing the number of code elements passed to the LLM. We also pruned the debug information by recording only variables that had changed and cropping long arrays and lists to further reduce the token size. Without such adjustments or modifications to Agentless—even when providing all the artifacts listed in this paper—its end-to-end performance remained below 15%. This demonstrates that directly applying a leading program repair framework designed for Python code to Java code does not yield competitive results.
Table 10. Resolution rates of frameworks using a GPT-4o model on the SWE-bench Lite Leaderboards
Tool LLM Resolved/Selected %Resolved
DEVLoRe GPT-4o-mini 46/207 22.2%
SWE-agent (Yang et al., 2024a) GPT-4o 55/300 18.3%
AppMap Navie (app, [n. d.]) GPT-4o 65/300 21.7%
Aider (aid, [n. d.]) GPT-4o+Claude-3 79/300 26.3%
Agentless (Xia et al., 2024a) GPT-4o 82/300 27.3%
MarsCode (Liu et al., 2024) Multiple LLMs(includes GPT-4o) 118/300 39.3%
Figure 12. A comparison between DEVLoRe with baselines on SWE-bench Lite
To compare with leading frameworks and evaluate DEVLoRe’s capability on Python code, we re-implemented our framework and tested it on SWE-bench Lite. The first challenge we encountered was that, even after filtering with MethodRecorder, the length of the related methods for one-third of the bugs in SWE-bench Lite still exceeded the LLM’s token limit. As a result, we were only able to provide related methods to the LLM for 207 out of 300 bugs in SWE-bench Lite. Additionally, we observed that Python’s debugging information is significantly more verbose than that of Java. For some bugs, the required patches involve modifying top-level constructs such as imports or global variables, for which debugging information cannot be extracted by executing the test cases. As shown in Table 10, among the 207 bugs with related methods information, DEVLoRe successfully fixed 46 (46/207=22.2%). We also provide a comparison with the SWE-bench Lite Leaderboards using the GPT-4o model. The results in Table 10 shows that DEVLoRe did not outperform Agentless or SWE-agent or MarsCode in the total number of fixed Python bugs. We suspect this is due to fundamental differences in coding practices between Java and Python, which result in related methods or debugging information being unavailable for a portion of the issues. Moreover, we compared the resolved issue IDs of DEVLoRe with those of baselines. The overlap of resolved issues among the baselines and DEVLoRe in Figure 12 shows that DEVLoRe was able to resolve 9 unique issues, despite resolving only 46 issues in total. Manual inspection of these 9 issues revealed that all of them involved the use of debugging information. This result suggests that the debugging information automatically extracted in our framework—among these baselines, only AppMap Naive utilizes runtime data when available in an interactive way—shows promise in addressing a subset of issues. Moreover, this experiment highlights that directly applying a leading framework from SWE-bench to Defects4J, or vice versa, may not yield state-of-the-art results. Generalizing the framework across different programming languages will be part of our future work.